OpenStudy (ttop0816):

What is the phase shift of y = csc2(x + pi/4)? -pi/2 -pi/4 <<< pi/4

OpenStudy (mathmale):

Compare your function to the general form:\[y=\csc(2(x+\frac{ \pi }{ 4 }))\rightarrow y=\csc(2x+\frac{ \pi }{ 2 })=\csc(bx+c).\] The period, in such case, is \[\frac{ 2\pi }{ b }\] and the phase shift is \[\frac{ -c }{ b }\] so, with c=pi/2 and b=2, I'd say the phase shift is \[\frac{ -c }{ b }=\frac{ -\frac{ \pi }{ 2 } }{ 2 }=\frac{ -\pi }{ 4 }.\]

OpenStudy (ttop0816):

@mathmale thanks! i also have a question about this

OpenStudy (ttop0816):

What is the phase shift of y = cos(3x - 3pi/4)? pi/4 3pi/4 9pi/4 4π if i do i divide the 3 from 3x to both cos ^ -3pi/4 or multiply them??

OpenStudy (mathmale):

Here I'd identify b as b=3 and c as c=-3pi/4. Thus, the phase shift would be -c/b. Try that.

OpenStudy (ttop0816):

how would i calculate -3pi/4 / 3?

OpenStudy (mathmale):

\[\frac{ -c }{ b }=\frac{ -(3/4)\pi }{ 3 }=\frac{ -\pi }{ 4 }\]

OpenStudy (ttop0816):

then the answer be pi/4 correct?!